# Enhancing Data Lakes with GraphAr: Efficient Graph Data Management with a Specialized Storage Scheme

### v3

@@ Abstract
Data lakes, increasingly adopted for their ability to store and analyze diverse types of data, commonly use columnar storage formats like Parquet and ORC for handling relational tables. However, these traditional setups fall short when it comes to efficiently managing graph data, particularly those conforming to the Labeled Property Graph (LPG) model. To address this gap, this paper introduces GraphAr, a specialized storage scheme designed to enhance existing data lakes for efficient graph data management. Leveraging the strengths of Parquet, GraphAr captures LPG semantics precisely and facilitates graph-specific operations such as neighbor retrieval and label filtering. Through innovative data organization, encoding, and decoding techniques, GraphAr dramatically improves performance. Our evaluations reveal that GraphAr outperforms conventional Parquet and Acero-based methods, achieving an average speedup of 3283×3283$\times$3283 × for neighbor retrieval, 6.0×6.0$\times$6.0 × for label filtering, and 29.5×29.5$\times$29.5 × for end-to-end workloads. These findings highlight GraphAr’s potential to extend the utility of data lakes by enabling efficient graph data management.
1 Introduction
@@
Data lakes have quickly become an essential infrastructure for organizations looking to store and analyze diverse datasets in their raw formats [64, 22, 28, 48, 47, 41, 53]. As centralized repositories, they offer unparalleled flexibility in accommodating a wide array of data types, from structured relational tables to unstructured logs and text files. Crucially, they serve as a cost-effective solution for archiving data at scale while still allowing for queries on archived or rarely accessed data. This dual utility makes them invaluable for both real-time analytics and long-term data management. Columnar storage formats like Parquet [5] and ORC [4] have become standard for storing tabular data in data lakes due to their robust compression and efficient query capabilities.
@@
In sync with these trends, graph data has become increasingly importance, especially for modeling complex relationships. Leading graph databases like Neo4j [13], TigerGraph [31], and JanusGraph [10] leverage the Labeled Property Graph (LPG) model [20, 23, 21] for this purpose. The recent ISO SQL:2023 standard includes a SQL/PGQ extension that not only facilitates querying property graphs but also enables the creation of property graph views from relational tables [19]. This groundbreaking inclusion highlights the growing convergence of relational and graph data models and emphasizes the need to integrate LPGs into data lakes. Consequently, LPGs are making their way into data lakes for multiple uses, from backups and archival storage for existing graph databases to natural extensions of transactional, log, and tabular data.
@@
Managing and analyzing LPGs in data lakes offers several significant benefits. Graph-specific queries are often more naturally articulated in languages like Cypher [35], Gremlin [55], GQL [34], or SQL/PGQ, providing an intuitive framework for conducting comprehensive analysis of entity relationships, facilitating the discovery of valuable insights. Data lakes also provide computational flexibility for running complex graph algorithms, enabling the exploration of intricate patterns. Additionally, they offer cost-effective storage solutions, allowing organizations to utilize more affordable and colder storage options without sacrificing query performance. Most notably, data lakes enable seamless querying across both graph and relational data, ushering in a holistic approach to data analytics.
Figure 1: A graph-related query within the data lake.
@@
Firstly, there is no standardized way to encapsulate an LPG within the existing data lake architecture. While columnar formats like Parquet and ORC excel at storing individual tables, they fall short in representing the complex relationships and semantics across these tables, which are inherent to LPGs.
Secondly, executing graph-specific operations, particularly neighbor retrieval and label filtering, can be highly inefficient in this setup. For example, neighbor retrieval might require multiple joins, significantly impacting performance. Label filtering, another essential operation, also introduces inefficiency due to the lack of native support in columnar formats.
GraphAr. To address these challenges, we introduce GraphAr, an efficient storage scheme for graph data in data lakes. GraphAr is designed to enhance the efficiency of data lakes while utilizing the capabilities of existing formats, with a specific focus on Parquet in this paper. GraphAr ensures seamless integration with existing tools and introduces innovative additions specifically tailored to handle LPGs.
Firstly, Parquet provides flexible and efficient support for various datatypes, including atomic types (e.g., bools and integers), and nested and/or repeated structures (e.g., arrays and maps). Using Parquet as fundamental building block, GraphAr further introduces standardized YAML files to represent the schema metadata for LPGs. This combination of data organization and metadata management enables the complete expression of LPG semantics, while ensuring compatibility with both data lakes and graph-related systems.
Secondly, GraphAr incorporates specialized optimization techniques to improve the performance of two critical graph operations: neighbor retrieval and label filtering, which are not inherently optimized in existing formats. GraphAr facilitates neighbor retrieval by organizing edges as sorted tables in Parquet to enable an efficient CSR/CSC-like representation, and leveraging Parquet’s delta encoding to reduce overhead in data storage and loading. GraphAr also introduces a novel decoding algorithm that utilizes BMI and SIMD instructions, along with a unique structure named PAC (page-aligned collections), to further accelerate the neighbor retrieval process. In addressing another critical aspect of LPGs, GraphAr adapts the run-length encoding (RLE) technique from Parquet and introduces a unique merge-based decoding algorithm. This tailored approach significantly improves label filtering performance, whether it involves simple or complex conditions.
@@
  • Development of specialized optimization techniques for enhancing performance in neighbor retrieval and label filtering operations (Sections 4 and 5).
  • Comprehensive performance evaluation of GraphAr compared to Parquet and Acero-based implementations, highlighting substantial speed gains: on average 3283×3283$\times$3283 × for neighbor retrieval, 6.0×6.0$\times$6.0 × for label filtering, and 29.5×29.5$\times$29.5 × for end-to-end workloads (Section 6).
2 Background and Key Challenges
In this section, we discuss the limitations of using tabular formats like Parquet and ORC in data lakes for LPGs, a common graph data model. We explore how these formats inadequately support LPG representation and efficient graph queries, laying the groundwork for the challenges that GraphAr tackles, outlined at the section’s end.
2.1 Tabular Formats in Data Lakes
Tabular data is key to data lakes, aiding efficient organization, analysis, and data extraction from large sets. Columnar formats like Parquet [5] and ORC [4] are popular due to their robust features. Unlike row-based formats such as CSV, they allow faster queries by enabling selective column reading, avoiding unnecessary data. Additionally, they offer diverse and efficient compression and encoding strategies, such as delta encoding to compress the variance between consecutive values, and run-length encoding to compress repetitive values. These techniques not only reduce storage needs but also enhance processing speeds. Another advantage of Parquet and ORC is predicate pushdown, which enhances query performance by moving filters closer to the storage layer, thereby reducing subsequent data reads.
The combination of selective column reading, efficient compression, and predicate pushdown positions Parquet and ORC as the go-to choices for managing tabular data in data lakes. Previous studies have demonstrated the importance of leveraging their capabilities for optimizing relational data management [33, 40, 22]. Recent research [45, 60, 61, 62, 63, 6] has also explored enhancements to tabular formats, utilizing CPU instruction sets like BMI and SIMD. In this paper, we will focus on Parquet, but the techniques discussed can be seamlessly adapted to other columnar formats such as ORC.
Figure 2: The internal structure of a Parquet file for a logical table with $C$ columns and $R$ row groups.
Parquet. Figure 2 illustrates the internal structure of a Parquet file. Structurally, a Parquet file represents a table, organized into row groups for logical segmentation. Within a row group, the data of a column is stored in a column chunk, which is guaranteed to be contiguous in the file. Column chunks are further divided into pages, which are the indivisible units for data compression and encoding. These pages, which can vary in type, are interleaved in a column chunk.
@@ 2.1 Tabular Formats in Data Lakes
2.2 Labeled Property Graphs
@@
Labeled Property Graphs (LPGs) [20, 23, 21] excel at representing complex relationships and semantics in a natural manner. Their flexible schema allows for accommodating the diverse and evolving nature of big data within repositories, making them integral to data lakes. LPG serves as the canonical data model in many graph systems [13, 36, 31] and graph query languages [59, 7, 35, 55, 34, 30], enabling queries and analytics to uncover valuable insights and patterns. Formally, an LPG is defined as $G=(V,E,T_{V},T_{E},P,L)$, where $V$ is a set of vertices, $E$ a set of edges, $T_{V}$ and $T_{E}$ the types of vertices and edges respectively, $P$ the properties, and $L$ the labels.
@@
For each vertex $v\in V$ and edge $e\in E$, they are associated with a type $t_{v}\in T_{V}$ and $t_{e}\in T_{E}$ respectively, and can have optional properties. A property $p\in P$ is specific to a vertex or an edge type, with a unique identifier within its type and a pre-defined datatype for its values. This implies that vertices or edges of the same type share the same set of properties.
@@
Furthermore, each label $l\in L$ has a unique identifier, usually a string. Each vertex type $t_{v}$ is linked with a set of candidate labels $L(t_{v})\subseteq L$, allowing each vertex of this type to be assigned zero or more labels. Labels hold significant importance in LPGs as they represent classifications and characteristics of entities, whereas properties serve as attributes to store additional information. Although edge types could technically also be labeled, in this paper, we focus solely on vertex labels, aligning with common graph query practice  [^1] Graph query languages like Cypher and Gremlin typically adhere to the convention that an edge can have only one classification, corresponding to the edge type in LPG model. Nevertheless, the strategies for vertex labels discussed in this paper can be seamlessly extended to support edge labels..
@@
While Parquet is highly effective for storing individual vertex and edge types $T_{V}$ and $T_{E}$ along with their associated properties due to the columnar structure and data compression capabilities, it falls short in capturing the interconnected schema essential for linking vertices with edges, e.g., to express the relationships across the three tables of Figure 3, which represent two vertex types and one edge type. This limitation is crucial for efficient graph traversal and pattern matching. Moreover, Parquet lacks native support for the multi-labeling capability of LPGs, resulting in a loss of complex semantics and relationships inherent to LPGs.
Figure 3: An example of querying LPGs on tabular formats.
2.3 Querying Labeled Property Graphs
@@
The core feature shared among existing graph query languages is the facility for pattern matching [30, 7]. This capability allows for in-depth analysis of the relationships between entities, uncovering valuable insights and patterns that may not be readily apparent in other data models. A graph pattern is defined as vertices and their connections through edges, filtered based on labels and property values [34]. Figure 3 illustrates the example workload mentioned in Figure 1, expressed in Cypher. And the steps for matching (a:Asian:Enrollee)-[e1:Diagnosed]->(d) are highlighted in the figure.
@@
When it comes to properties, a viable approach is to use native tabular data. This leverages the storage efficiency of existing formats, while optimizing property-related operations in graph queries. However, this approach struggles with two crucial aspects of pattern matching.
@@
Firstly, tabular formats lack native support for representing graph topology, making it difficult to efficiently fetch the neighboring vertices and edges for a given vertex. A common workaround is to store edge endpoints as properties and use the join operations across multiple tables to retrieve neighbors, as shown in Figure 3. However, this approach is often inefficient due to the computational overhead of multiple joins.
@@
Secondly, LPGs allow vertices to have multiple labels, offering a flexible and expressive way to describe entities. Label filtering is an essential feature in graph queries, to enable the selection of specific subsets of vertices, making it a core element in all graph query languages [34, 7, 55, 59]. Existing formats do not natively accommodate this flexibility and do not provide a foundation for label-based optimizations. Encoding labels as ordinary properties and performing filtering by string matching, as seen in the initial step of Figure 3, limits the expressive power of vertex representation and hampers efficient label handling, a feature unique to graph queries.
@@
2.4 Key Challenges Addressed by GraphAr

### v4

# GraphAr: An Efficient Storage Scheme for Graph Data in Data Lakes

@@ Abstract.
Data lakes, increasingly adopted for their ability to store and analyze diverse types of data, commonly use columnar storage formats like Parquet and ORC for handling relational tables. However, these traditional setups fall short when it comes to efficiently managing graph data, particularly those conforming to the Labeled Property Graph (LPG) model. To address this gap, this paper introduces GraphAr, a specialized storage scheme designed to enhance existing data lakes for efficient graph data management. Leveraging the strengths of Parquet, GraphAr captures LPG semantics precisely and facilitates graph-specific operations such as neighbor retrieval and label filtering. Through innovative data organization, encoding, and decoding techniques, GraphAr dramatically improves performance. Our evaluations reveal that GraphAr outperforms conventional Parquet and Acero-based methods, achieving an average speedup of 4452×4452$\times$4452 × for neighbor retrieval, 14.8×14.8$\times$14.8 × for label filtering, and 29.5×29.5$\times$29.5 × for end-to-end workloads. These findings highlight GraphAr’s potential to extend the utility of data lakes by enabling efficient graph data management.
PVLDB Reference Format:
Xue Li, Weibin Zeng, Zhibin Wang, Diwen Zhu, Jingbo Xu, Wenyuan Yu, Jingren Zhou. GraphAr: An Efficient Storage Scheme for Graph Data in Data Lakes. PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/apache/incubator-graphar/tree/research.
1. Introduction
@@
Data lakes have quickly become an essential infrastructure for organizations looking to store and analyze diverse datasets in their raw formats (Zaharia et al., 2021; Armbrust et al., 2020; Dageville et al., 2016; Nargesian et al., 2019; Miloslavskaya and Tolstoy, 2016; Khine and Wang, 2018; Ramakrishnan et al., 2017). As centralized repositories, they offer unparalleled flexibility in accommodating a wide array of data types, from structured relational tables to unstructured logs and text files. Crucially, they serve as a cost-effective solution for archiving data at scale while still allowing for queries on archived or rarely accessed data. This dual utility makes them invaluable for both real-time analytics and long-term data management. Columnar storage formats like Parquet (par, 2024) and ORC (orc, 2024) have become standard for storing tabular data in data lakes due to their robust compression and efficient query capabilities.
@@
In sync with these trends, graph data has become increasingly importance, especially for modeling complex relationships. Leading graph-related systems like Neo4j (neo, 2024b), TigerGraph (Deutsch et al., 2019), JanusGraph (jan, 2024) and GraphScope (Fan et al., 2021) leverage the Labeled Property Graph (LPG) model (Angles et al., 2023; Baken, 2020; Anikin et al., 2019) for this purpose. The recent ISO SQL:2023 standard includes a SQL/PGQ extension that not only facilitates querying LPGs but also enables the creation of LPG views from relational tables (SQL, 2023). This groundbreaking inclusion highlights the growing convergence of relational and graph data models and emphasizes the need to integrate LPGs into data lakes. Consequently, LPGs are making their way into data lakes for multiple uses, from backups and archives for existing graph databases to natural extensions of transactional, log, and tabular data.
@@
Managing and analyzing LPGs in data lakes offers several significant benefits. Graph-specific queries are often more naturally articulated in languages like Cypher (Francis et al., 2018), Gremlin (Rodriguez, 2015), GQL (Francis et al., 2023), or SQL/PGQ, providing an intuitive framework for conducting comprehensive analysis of entity relationships, facilitating the discovery of valuable insights. Data lakes also provide computational flexibility for running complex graph algorithms, enabling the exploration of intricate patterns. Additionally, they offer cost-effective storage solutions, allowing organizations to utilize more affordable and colder storage options without sacrificing query performance. Most notably, data lakes enable seamless querying across both graph and relational data, ushering in a holistic approach to data analytics.
Figure 1. A graph-related query within the data lake.
@@
Challenge 1: There is no standardized way to encapsulate an LPG within the existing data lake architecture. While columnar formats like Parquet and ORC excel at storing individual tables, they fall short in representing the complex relationships and semantics across these tables, which are inherent to LPGs.
Challenge 2: Graph-specific operations can be highly inefficient in this setup. The foundational operation, neighbor retrieval, might require multiple joins, significantly impacting performance.
Challenge 3: Label filtering, another essential graph-specific operation, also introduces inefficiency due to the lack of native support in columnar formats.
GraphAr. To address these challenges, we introduce GraphAr  [^1] Apache GraphAr is an effort undergoing incubation at the Apache Software Foundation (ASF), sponsored by the Apache Incubator., an efficient storage scheme for graph data in data lakes. It is designed to enhance the efficiency of data lakes utilizing the capabilities of existing formats, with a specific focus on Parquet in this paper. GraphAr ensures seamless integration with existing tools and introduces innovative additions specifically tailored to handle LPGs.
Firstly, Parquet provides flexible and efficient support for various datatypes, including atomic types (e.g., bools and integers), and nested and/or repeated structures (e.g., arrays and maps). Leveraging Parquet as its fundamental building block, GraphAr further introduces standardized YAML files to represent the schema metadata for LPGs, alongside a hierarchical data layout to store the data. This innovative combination of data organization with metadata management enables the complete expression of LPG semantics, while ensuring compatibility with both data lake ecosystems and existing graph-related systems, addressing Challenge 1.
GraphAr incorporates specialized optimization techniques to improve the performance of critical graph operations, which are not inherently optimized in existing formats. To address Challenge 2, GraphAr facilitates neighbor retrieval by organizing edges as sorted tables in Parquet to enable an efficient CSR (Compressed Sparse Row) or CSC (Compressed Sparse Column)-like representation, and leveraging Parquet’s delta encoding to reduce overhead in data storage and loading. GraphAr also introduces an innovative decoding algorithm that utilizes instruction sets such as BMI (Bit Manipulation Instructions) and SIMD (Single Instruction, Multiple Data), along with a unique structure named PAC (Page-Aligned Collections), to further accelerate the neighbor retrieval process.
In addressing Challenge 3 of label filtering, GraphAr adapts the RLE (Run-Length Encoding) technique from Parquet and introduces a novel interval-based decoding algorithm. Through integrating proven methods (CSR/CSC, delta encoding, RLE) with novel decoding algorithms, GraphAr delivers a comprehensive and efficient solution for optimizing LPG-specific operations.
@@
  • Development of specialized optimization techniques for enhancing performance in neighbor retrieval and label filtering operations (Sections 4 and 5). These are built upon Parquet’s advanced encoding features and are complemented by two innovative and efficient decoding algorithms.
  • Comprehensive performance evaluation of GraphAr compared to Parquet and Acero-based implementations, highlighting substantial speed gains: on average 4452×4452$\times$4452 × for neighbor retrieval, 14.8×14.8$\times$14.8 × for label filtering, and 29.5×29.5$\times$29.5 × for end-to-end workloads. And the potential for integrating GraphAr into existing graph systems (Section 6).
2. Background and Key Challenges
In this section, we discuss the limitations of using tabular file formats like Parquet and ORC in data lakes for LPGs, a common graph data model. We explore how these formats inadequately support LPG representation and efficient graph queries, laying the groundwork for the challenges that GraphAr tackles.
2.1. Tabular Formats in Data Lakes
Tabular data is key to data lakes, aiding efficient organization, analysis, and data extraction from large sets. Columnar formats like Parquet (par, 2024) and ORC (orc, 2024) are popular due to their robust features. Unlike row-based formats such as CSV, they allow faster queries by enabling selective column reading, avoiding unnecessary data. Additionally, they offer diverse and efficient compression and encoding strategies, such as delta encoding to compress the variance between consecutive values, and run-length encoding to compress repetitive values. These techniques not only reduce storage needs but also enhance processing speeds. Another advantage of Parquet and ORC is predicate pushdown, which enhances query performance by moving filters closer to the storage layer, thus minimizing reads.
The combination of selective column reading, efficient compression, and predicate pushdown positions Parquet and ORC as the go-to choices for managing tabular data in data lakes. Previous studies have demonstrated the importance of leveraging their capabilities for optimizing relational data management (Floratou et al., 2014; Ivanov and Pergolesi, 2020; Armbrust et al., 2020). Recent research (Li et al., 2023; Willhalm et al., 2013, 2009; Yang et al., 2021; Yu et al., 2020; AWS, 2024) has also explored enhancements to tabular formats, utilizing CPU instruction sets like BMI and SIMD. In this paper, we will focus on Parquet, but the techniques discussed can be seamlessly adapted to other columnar formats such as ORC.
Figure 2. The internal structure of a Parquet file for a logical table with $C$ columns and $R$ row groups.
Parquet. Figure 2 illustrates the internal structure of a Parquet file. Structurally, a Parquet file represents a table, organized into row groups for logical segmentation. Within a row group, the data of a column is stored in a column chunk, which is guaranteed to be contiguous in the file. Column chunks are further divided into pages, the indivisible units for compression and encoding. These pages, which can vary in type, are interleaved in a column chunk.
@@ 2.1. Tabular Formats in Data Lakes
2.2. Labeled Property Graphs
@@
Labeled Property Graphs (LPGs) (Angles et al., 2023; Baken, 2020; Anikin et al., 2019) excel at representing complex relationships and semantics in a natural manner. Their flexible schema allows for accommodating the diverse and evolving nature of big data within repositories, making them integral to data lakes. LPG serves as the canonical data model in many graph systems (neo, 2024b; Gonzalez et al., 2014; Deutsch et al., 2019) and graph query languages (van Rest et al., 2016; cyp, 2024; Francis et al., 2018; Rodriguez, 2015; Francis et al., 2023; Deutsch et al., 2022), enabling queries and analytics to uncover valuable insights and patterns. Formally, an LPG is defined as $G=(V,E,T_{V},T_{E},P,L)$, where $V$ is a set of vertices, $E$ a set of edges, $T_{V}$ and $T_{E}$ the types of vertices and edges respectively, $P$ the properties, and $L$ the labels.
@@
For each vertex $v\in V$ and edge $e\in E$, they are associated with a type $t_{v}\in T_{V}$ and $t_{e}\in T_{E}$ respectively, and can have optional properties. A property $p\in P$ is specific to a vertex or an edge type, with a unique identifier within its type and a pre-defined datatype for its values. This implies that vertices or edges of the same type share the same set of properties, and can be stored in a tabular format, as shown in Figure 3(a).
@@
Furthermore, each label $l\in L$ has a unique identifier, usually a string. Each vertex type $t_{v}$ is linked with a set of candidate labels $L(t_{v})\subseteq L$, allowing each vertex of this type to be assigned zero or more labels. Labels hold significant importance in LPGs as they represent classifications and characteristics of entities, while properties serve as attributes to store additional information. LPGs allow vertices to have multiple labels, offering a flexible and expressive way to describe entities. For instance, Figure 3(b) illustrates an inheritance hierarchy of labels within the Person vertices, where a vertex can be labeled as both Asian and Enrollee. Although edge types could technically also be labeled, we focus solely on vertex labels in this paper, aligning with common graph query practice [^2] Graph query languages like Cypher and Gremlin typically adhere to the convention that an edge can have only one classification, corresponding to the edge type in LPG model. Nevertheless, the strategies for vertex labels discussed in this paper can be seamlessly extended to support edge labels..
@@
While Parquet is highly effective for storing individual vertex and edge types $T_{V}$ and $T_{E}$ along with their associated properties due to the columnar structure and data compression capabilities, it falls short in capturing the interconnected schema essential for linking vertices with edges, e.g., to express the relationships across the three tables of Figure 3(a), which represent two vertex types and one edge type. This limitation is crucial for efficient graph traversal and pattern matching. Moreover, Parquet lacks native support for the multi-labeling capability of LPGs, resulting in a loss of complex semantics and relationships inherent to LPGs.
(a) Execution workflow of the query on tabular formats.
(b) Inherent multi-labeling characteristics inside Person vertices.
Figure 3. An example of querying LPGs on tabular formats.
2.3. Querying Labeled Property Graphs
@@
The core feature shared among graph query languages is the facility for pattern matching (Deutsch et al., 2022; cyp, 2024). This capability allows for in-depth analysis of the relationships between entities, uncovering valuable insights and patterns that may not be readily apparent in other data models. Given the fact that a LPG consists of topology, labels, and properties, a graph pattern is then defined as vertices and their connections through edges, filtered based on labels and property values (Francis et al., 2023). Figure 3(a) illustrates the example workload mentioned in Figure 1, expressed in Cypher. And the steps for matching (a:Asian:Enrollee)-[e1:Diagnosed]->(d) are highlighted in the figure.
@@
When it comes to properties, a viable approach is to use native tabular data, leveraging existing formats for efficient storage and property-related operations. However, this approach struggles with two crucial aspects of pattern matching.
@@
Firstly, tabular formats lack native support for representing graph topology, making it difficult to efficiently fetch the neighboring vertices and edges for a given vertex. A common workaround is to store edge endpoints as properties and use the join operations across multiple tables to retrieve neighbors, as shown in Figure 3(a). However, this approach is often inefficient due to the computational overhead of multiple joins.
@@
Secondly, label filtering is a unique feature in graph queries, to enable the selection of specific subsets of vertices, making it an essential element in all graph query languages (Francis et al., 2023; cyp, 2024; Rodriguez, 2015; van Rest et al., 2016). Existing formats do not natively accommodate this flexibility and do not provide a foundation for label-based optimizations. Encoding labels as ordinary properties and performing filtering by string matching, as seen in the initial step of Figure 3(a), limits the expressive power of vertex representation and hampers efficient label handling.
@@
2.4. Key Challenges Addressed by GraphAr